### v2

@@ 3.4.1 Empirical Evaluation
In contrast, in the spML scenario, assuming the asserter controls a fraction $r=10\%$ of the network, by Proposition 1, the probability of triggering the challenge mechanism is $0.735\%$. This translates to only $0.735\%$ additional computational overhead in spML, enabling us to completely avoid fraud and eliminate the need for fraud proof procedures, if all nodes are rational. Hence, spML is clearly the superior choice.
@@ 4 Conclusions and Future Extensions
In this study, we have introduced a robust PoSP protocol applicable across all kinds of scenarios, such as decentralized AI inference network, distinguished by its exceptional security features. Central to its design is the implementation of a unique Nash Equilibrium in pure strategies, ensuring that all rational participants within the network adhere to honest behavior. This protocol demonstrates superior performance when compared to existing methodologies across various applications.
Looking ahead, further exploration into the application of the PoSP protocol within Layer 2 architectures holds promise, particularly by employing our method of sampling multiple nodes to recompute results. This approach can lead to a unique Nash Equilibrium in pure strategies, where every participant acts honestly, directly addressing and potentially solving the concerns highlighted in [14]. Following this, there is significant potential for utilizing the PoSP protocol in the development of verification mechanisms within Actively Validated Services (AVS), enhancing security and reliability in EigenLayer. This exploration could lead to innovative applications that leverage the strengths of PoSP in ensuring robust, scalable, and secure systems.
@@
First, we consider the expected payoff of the asserter if he does not commit fraud. If the asserter does not commit fraud, the expected payoff is
$(1-p)(R_{A}-C)+p\sum_{i=0}^{n}\binom{n}{i}r^{i}(1-r)^{n-i}\left(i\frac{R_{V}}{n}+R_{A}-C\right).$
This is because when challenge mechanism is triggered, if $i$ out of $n$ selected validators collude with the asserter,the collusions can have the payoff $iR_{v}/n+R_{A}-C$, because the colluders are able to obtain the outcome of $f(x)$ directly from the asserter instead of recalculating it themselves.
Then, we consider the expected payoff of the asserter if he commits fraud. If the challenge mechanism is not triggered, the asserter is capable of compromising the integrity of the system and get a payoff of $U_{1}$. However, if the challenge mechanism is triggered, by Property 1, only when all of the $n$ selected validators collude with the asserter, the fraud will not be detected, and the asserter can earn $U_{2}$. If $1\leq i<n$ out of $n$ selected validators collude with the asserter, their optimal strategy is to act honestly and report the fraud to get the reward for validation rather than be slashed. Thus, the expected payoff is
$(1-p)U_{1}+pr^{n}U_{2}+p\sum_{i=0}^{n-1}\binom{n}{i}r^{i}(1-r)^{n-i}\left(i\frac{R_{V}}{n}-C-S\right).$
Hence, the system having a unique Nash Equilibrium in pure strategies is equivalent to the scenario where the asserter can obtain a greater profit if he does not commit fraud, which is
$(1-p)(R_{A}-c)+p\sum_{i=0}^{n}\binom{n}{i}r^{i}(1-r)^{n-i}\left(i\frac{R_{V}}{n}+R_{A}-C\right)$
$>$ $(1-p)U_{1}+pr^{n}U_{2}+p\sum_{i=0}^{n-1}\binom{n}{i}r^{i}(1-r)^{n-i}\left(i\frac{R_{V}}{n}-C-S\right).$
@@
$R_{A}+pS-(1-p)U_{1}-C+(1-r^{n}-(1-r)^{n})pC>pr^{n}(U_{2}+S-R_{V}),$

### v3

@@ 3.7.1 Empirical Evaluation
In contrast, in the spML scenario, assuming the fraction of the Byzantine nodes in the network $r=10\%$, by Proposition 1, the probability of triggering the challenge mechanism is $0.736\%$. This translates to only $0.736\%$ additional computational overhead in spML, enabling us to completely avoid fraud and eliminate the need for fraud proof procedures, if all nodes are rational. Hence, spML is the superior choice.
@@ 4 Conclusions and Future Extensions
In this study, we introduced the PoSP protocol and demonstrated a key application to decentralized AI inference network. Central to its design is the implementation of a unique Nash Equilibrium in pure strategies, ensuring that all rational participants within the network adhere to correct outputs, under certain assumptions. This protocol demonstrates superior performance when compared to zero-knowledge based protocols.
Looking ahead, further exploration into the application of the PoSP protocol within Layer 2 architectures holds promise, particularly by employing our method of sampling multiple nodes to recompute results. This approach can lead to a unique Nash Equilibrium in pure strategies, where every participant acts honestly, directly addressing and potentially solving the concerns highlighted in [25]. Additionally, there is significant potential for applying PoSP as a verification mechanism within Actively Validated Services (AVS) in restaking protocols. Recent research highlights that fragmented validator sets can compromise security unless stakes are dynamically rebalanced [27]. Industry developments underscore a transition towards autonomously verifiable services, reinforcing the necessity for efficient, lightweight verification mechanisms. This exploration could lead to innovative applications that leverage the strengths of PoSP in ensuring robust, efficient, and reliable systems.
Acknowledgements.
We are thankful to Ari Juels, Zhongjing Wei, Zhe Ye, Jianzhu Yao, Chenghan Zhou, Yuchen Jin, and Dahlia Malkhi for helpful comments and conversations.
@@ Acknowledgements.
First, we consider the expected payoff of the asserter if his result is correct. If the asserter’s result is correct, all the validators, whether Byzantine or not, have a dominant strategy to output the correct result. Suppose $k$ is the number of nodes controlled by the asserter that are selected as validators when the challenge mechanism is triggered. If the asserter does not commit fraud, the expected payoff for the asserter in this round is at least
$(1-p)(R_{A}-C)+p\left(\frac{R_{V}}{n}\mathbb{E}\left[k\right]+R_{A}-C\right).$
Then, we consider the expected payoff of the asserter if he commits fraud and outputs the incorrect result. If the challenge mechanism is not triggered, the asserter can get a payoff of $U_{1}$. However, if the challenge mechanism is triggered, fraud might go undetected and the asserter could earn $U_{2}$ only if all of the $n$ selected validators are Byzantine and submit the same result as the asserter. If $1\leq i<n$ out of $n$ selected validators are Byzantine or even collude with the asserter, their optimal strategy remains to act honestly and report the fraud in order to receive the validation reward. This is because the arbitration process will inevitably be triggered by the presence of at least one honest validator, making any dishonest action be penalized rather than rewarded.
Suppose $\rho\leq r$ is the fraction of Byzantine nodes that potentially submit the same result as the asserter out of all the nodes in the network; this includes nodes that may be controlled by the asserter as well. The expected payoff of the asserter is at most
$(1-p)U_{1}+p\rho^{n}U_{2}+p\sum_{i=0}^{n-1}\binom{n}{i}\rho^{i}(1-\rho)^{n-i}\left(\frac{R_{V}}{n}\mathbb{E}\left[k\middle|m=i\right]-S\right),$
where $m$ is the number of Byzantine nodes that potentially submit the same result as the asserter.
Hence, the system will have a unique Nash Equilibrium in pure strategies when the Byzantine asserter can obtain a greater profit if it does not commit fraud, i.e., if
$(1-p)(R_{A}-C)+p\left(\frac{R_{V}}{n}\mathbb{E}\left[k\right]+R_{A}-C\right)>(1-p)U_{1}+p\rho^{n}U_{2}+p\sum_{i=0}^{n-1}\binom{n}{i}\rho^{i}(1-\rho)^{n-i}\left(\frac{R_{V}}{n}\mathbb{E}\left[k\middle|m=i\right]-S\right).$
@@
$R_{A}+pS-(1-p)U_{1}-C>p\rho^{n}\left(U_{2}+S-\frac{R_{V}}{n}\mathbb{E}\left[k|m=n\right]\right).$
Since we always have $\mathbb{E}\left[k|m=n\right]\geq 0$ and $\rho\leq r$, our system will have a unique Nash Equilibrium in pure strategies, if
$R_{A}+pS-(1-p)U_{1}-C>pr^{n}\left(U_{2}+S\right),$